# CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction

### v1

@@ 4.1 Experimental Settings
Trade-off factors.
We employ a cross-evaluation approach to determine two optimal configurations for trade-off factors applicable at the corpus and sentence levels, respectively. At the corpus level, we assign the factors as . Conversely, at the sentence level, these are adjusted to . Additionally, CLEME2.0 metrics named with “sim” are based on similarity-based edit weighting, and we leave the analysis of LLM-based edit weighting to Section 5.2.
Evaluation Assumptions.
Ye et al. (2023c) propose evaluating systems based on one of two assumptions, namely correction dependence and correction independence. In short, the correction independence assumption offers a more relaxed edit matching process, implying that systems are more inclined to yield higher scores when multiple references are accessible. Inspired by this work, CLEME2.0 also supports both assumptions, and we will study their impact on our method.
4.2 Results of GJG15 Ranking
For corpus-level, CLEME2.0-sim-ind achieves the highest average correlations, closely followed by CLEME2.0-sim-dep. CLEME2.0-ind and CLEME2.0-dep can also gain comparable correlations with other metrics, in spite of the fact that they do not utilize any edit weighting techniques. On the other hand, sentence-level metrics exhibit a similar pattern. SentCLEME2.0-sim-dep and SentCLEME2.0-sim-ind achieve the highest Pearson and Spearson correlations, respectively. These results significantly demonstrate the effectiveness and robustness of our proposed method across different settings.
This observation aligns with recent studies Gong et al. (2022); Ye et al. (2023c). This is because system-level rankings treat each sample equally, which is consistent with the evaluation process of sentence-level metrics. In contrast, corpus-level metrics allow samples with more edits to significantly influence the results. SentPT-M2 exhibits the best performance on CoNLL-2014, but its results on BN-10GEC, E-Minimal, and NE-Fluency are inferior compared to our method.
In general, our method aligns more consistently with human judgments than existing mainstream metrics. Particularly, most weighted outcomes outshine the unweighted ones, attributable to the incorporation of semantic considerations. However, on E-Minimal and NE-Minimal, the unweighted and weighted results exhibit comparability. We conjecture that this could be due to the annotations in these datasets being minimal yet decisive, reducing the possibility of generating diverse weights.
4.3 Results of SEEDA Ranking
We conduct a supplementary experiment on the SEEDA-Sentence and SEEDA-Edit datasets, comparing our method with a wide range of GEC metrics. Table 2 demonstrates again that our approach obtains the optimal results on both datasets. Kobayashi et al. (2024b) claim that the correlations of most metrics tend to decline when shifting from classical to neural systems in evaluation. This suggests that traditional metrics may struggle when assessing more extensively edited and fluent corrections generated by neural systems. However, our method is still able to address these challenges and obtain even better results. The results on SEEDA-Edit surpass those on SEEDA-Sentence due to the finer granularity of SEED-Edit, which is more consistent with the functioning of CLEME2.0.
@@ 4.3 Results of SEEDA Ranking
| Dataset | Corpus-EW | Corpus-TS | Sentence-EW | Sentence-TS | ||||
| CoNLL-2014 | 0.697 | 0.659 | 0.759 | 0.720 | 0.626 | 0.654 | 0.696 | 0.698 |
| BN-10GEC | 0.732 | 0.764 | 0.796 | 0.813 | 0.638 | 0.637 | 0.708 | 0.698 |
| E-Minimal | 0.709 | 0.786 | 0.779 | 0.819 | 0.642 | 0.692 | 0.715 | 0.747 |
| E-Fluency | 0.760 | 0.786 | 0.831 | 0.841 | 0.642 | 0.665 | 0.720 | 0.714 |
| NE-Minimal | 0.777 | 0.823 | 0.839 | 0.861 | 0.654 | 0.747 | 0.723 | 0.791 |
| NE-Fluency | 0.823 | 0.692 | 0.849 | 0.709 | 0.664 | 0.791 | 0.742 | 0.830 |
| Source | When we are | diagonosed out | with certain genetic | disease | , are we suppose to disclose | this result | to | our | relatives ? |
| Ref. | When we are | diagnosed | with certain genetic | diseases | , are we suppose to disclose | this result | to | our | relatives ? |
| Hyp. | When we are | diagnosed out (0.056) | with certain genetic | diseases (0.006) | , are we suppose to disclose | the results (0.019) | to | their (0.021) | relatives ? |
| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | |
| Ref | Does one | who | suffers | from this disease keep it a secret | or inform | their relatives ? |
| Hyp. | Do one (0.028) | who | suffer (0.011) | from this disease keep it a secret | to inform (0.094) | their relatives ? |
5 Analysis
5.1 Ablation Studies
Isolated effect of each disentangled score.
We perform ablation experiments on (Sent)CLEME2.0-dep to observe how each disentangled score performs. Since a system exhibiting lower error-correction, under-correction, and over-correction is more desirable, these scores are subtracted from 1. The results are presented in Table 3. Both hit-correction and under-correction scores display moderate correlations. Over-correction scores exhibit slight positive correlations at the corpus-level, but negligible negative correlations at the sentence level. Interestingly, error-correction scores manifest negative correlations. However, this does not imply that error-correction scores have no influence on the comprehensive score. In fact, the trade-off factor of error-correction scores is relatively significant. It is hypothesized that evaluations based solely on error-correction scores could unduly encourage systems that produce only highly confident edits, resulting in potential evaluation bias.
Average correlations.

### v2

@@ 4.2 Results of GJG15 Ranking
4.3 Results of SEEDA Ranking
We carry out an additional experiment on the SEEDA-Sentence and SEEDA-Edit datasets, where we compare our method against various GEC metrics. As presented in Table 2, our approach consistently achieves the best outcomes across both datasets. According to Kobayashi et al. (2024b), the correlations of most metrics tend to decrease when transitioning from classical to neural evaluation systems. This implies that conventional metrics might face difficulties in evaluating the more extensively edited or fluent corrections produced by state-of-the-art neural GEC systems. Nevertheless, our method effectively tackles these challenges, delivering even improved performance for all indicators. The results for SEEDA-Edit exceed those for SEEDA-Sentence, due to the greater detail in SEEDA-Edit, aligning more closely with the operation of CLEME2.0.
| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | |
| Reference | Does one | who | suffers | from this disease keep it a secret | or inform | their relatives ? |
| Hypothesis | Do one (0.028) | who | suffer (0.011) | from this disease keep it a secret | to inform (0.094) | their relatives ? |
| Source | When we are | diagonosed out | with certain genetic | disease | , should we disclose | this result | to | our | relatives ? |
| Ref. | When we are | diagnosed | with certain genetic | diseases | , should we disclose | this result | to | our | relatives ? |
| Hyp. | When we are | diagnosed out (0.056) | with certain genetic | diseases (0.006) | , should we disclose | the results (0.019) | to | their (0.021) | relatives ? |
| Dataset | Corpus-EW | Corpus-TS | Sentence-EW | Sentence-TS | ||||
| CoNLL-2014 | 0.697 | 0.659 | 0.759 | 0.720 | 0.626 | 0.654 | 0.696 | 0.698 |
| BN-10GEC | 0.732 | 0.764 | 0.796 | 0.813 | 0.638 | 0.637 | 0.708 | 0.698 |
| E-Minimal | 0.709 | 0.786 | 0.779 | 0.819 | 0.642 | 0.692 | 0.715 | 0.747 |
| E-Fluency | 0.760 | 0.786 | 0.831 | 0.841 | 0.642 | 0.665 | 0.720 | 0.714 |
| NE-Minimal | 0.777 | 0.823 | 0.839 | 0.861 | 0.654 | 0.747 | 0.723 | 0.791 |
| NE-Fluency | 0.823 | 0.692 | 0.849 | 0.709 | 0.664 | 0.791 | 0.742 | 0.830 |
4.4 Results of LLM-based Weighting
Table 4 presents the outcomes of LLM-based weighting, noting that its effectiveness is less favorable than similarity-based weighting. A likely reason is the coarse grading method of LLMs, which allocates edit weights from 1 to 5, unlike the finer continuous scale . Although Kobayashi et al. (2024a) argue that LLMs serve as effective evaluators for GEC, their research pertains to huge closed-source LLMs (GPT-4 and GPT-3.5) and involves specific prompt engineering. They also identify the importance of the LLM scale since GPT-3.5 may even obtain negative correlations with human judgments. In contrast, we employ a more straightforward approach with open-source LLama-2-7B.
5 Analysis
5.1 Case Study
Table 3 demonstrates instances of CLEME2.0. In the first set, Chunks 3 and 5 are FPne edits contributing to the wrong-correction score, with a higher edit weight of Chunk 5 than Chunk 3 since Chunk 5 introduces an error that entirely alters the sentence’s meaning. In the second set, Chunk 2 obtains the highest edit weight of 0.056, underscoring its substantial influence on the evaluation. Despite the correct modification of “diagnosed”, the misuse of “out” remains, keeping the correction wrong. Chunk 4 illustrates a singular-to-plural correction in the source sentence, with a low weight indicating a minor impact. Chunks 6 and 8 showcase over-corrections. Chunk 6 leaves the original meaning unchanged, whereas Chunk 8 introduces a significant error by misusing a personal pronoun.
The cases highlight the effectiveness of the weighting technique. Otherwise, all edits are given equal weight, failing to distinguish hypothesis edits with varying correction levels. We display the cases of LLM-based weighting in Appendix C.2.
5.2 Ablation Study
@@
We conduct ablation studies on (Sent)CLEME2.0-dep to analyze the performance of individual disentangled scores. A preferable system has reduced wrong-correction, under-correction, and over-correction scores, so we report corrections between 1- with human judgments where is one of the scores. The outcomes are detailed in Table 5. Hit-correction and under-correction show moderate correlations. Over-correction scores have small positive correlations at the corpus level, with minimal negative correlations at the sentence level. Notably, wrong-correction scores display negative correlations, but this does not mean they do not impact the overall score. In reality, the trade-off factor for wrong-correction scores is relatively substantial. The hypothesis is that focusing evaluations only on wrong-correction scores might prefer systems that make only highly confident edits, potentially leading to assessment bias.
Additionally, we utilize the similarity-weighting approach on CLEME to evaluate its efficacy, with the outcomes detailed in Appendix C.3. To examine our method on a broad scale, we also provide the average correlations obtained from a comprehensive analysis of all potential parameter settings. The results are found in Appendix C.4.
| Metric | Time (Seconds) |
| ERRANT | 33.4 |
| GLEU | 21.5 |
| CLEME-dep | 54.1 |
| CLEME-ind | 54.1 |
| (Sent)CLEME2.0-dep | 54.1 |
| (Sent)CLEME2.0-ind | 54.1 |
| (Sent)CLEME2.0-sim-ind | 88.4 |
| (Sent)CLEME2.0-sim-ind | 87.6 |
5.3 Efficiency
This section provides a comparative analysis of the efficiency of our methods against other prevailing metrics. The experiments were executed on a GPU 3090 within the CoNLL-2014 framework, with the evaluation times of the AMU system reported. Our observations are as follows: (1) For ERRANT, the primary time expenditure is associated with edit extraction, lasting 33.4 seconds. (2) CLEME and CLEME2.0 primarily incur time costs from edit extraction at 33.4 seconds and chunk partitioning at 20.7 seconds. (3) For CLEME2.0-sim, the most significant time costs are assignable to edit extraction (33.4 s), chunk partitioning (20.7 s), and edit weighting (34.3 s). PT-M2 exhibits the slowest runtime when replicating existing mainstream methods, with its evaluation process taking several hours; thus, we did not report a precise runtime due to the time constraints. Some technical solutions can mitigate the runtime when evaluating a system using these metrics concurrently. For instance, when assessing a system with ERRANT, CLEME, and CLEME2.0, the minimum cumulative duration is calculated as 33.4 seconds for edit extraction, 20.7 seconds for chunk partitioning, and 34.3 seconds for edit weighting, totaling 88.4 seconds.
6 Conclusion
This paper introduces CLEME2.0, an interpretable evaluation metric for GEC that effectively highlights four key aspects of systems. By incorporating edit weighting techniques, we overcome the challenges traditional reference-based metrics face in recognizing semantic subtleties. Extensive experiments and analyses confirm the effectiveness and robustness of our method. We anticipate that CLEME2.0 will offer a valuable perspective in the GEC community.
Limitation
Limitation in languages and datasets.
While CLEME2.0 is adaptable to various languages, its efficiency beyond English remains unverified. Additionally, the reference sets employed in our experiments stem from the CoNLL-2014 shared task, which involves a second language dataset. To confirm the robustness of our methods, it’s necessary to conduct further experiments using evaluation datasets that cover a range of languages and text domains. Finally, we highly encourage the creation of new GEC evaluation datasets to foster progress.
Lack of further human evaluation for interpretability.
The experiments discussed in the paper are primarily concerned with assessing the correlation between automatic metrics and human judgments. However, they fall short of providing a thorough analysis of the method’s interpretability. Although we showcase the strong correlation performance of CLEME2.0, its interpretability is still unverified. In future work, we will conduct human evaluation experiments to showcase the interpretability of our method.
Ethics Statement
Acknowledgements
This research is supported by National Natural Science Foundation of China (Grant No. 62276154), Research Center for Computer Network (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No. 2023A1515012914 and 440300241033100801770), Basic Research Fund of Shenzhen City (Grant No. JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006), the Major Key Project of PCL (NO. PCL2024A08).
Appendix A LLM-based Edit Weighting
Because of the powerful semantic comprehension abilities of LLMs Qin et al. (2024a); Tan et al. (2024); Ye et al. (2025a); Yu et al. (2024a); Tang et al. (2025); Yan et al. (2025); Li et al. (2024e, 2022a); Du et al. (2024); Huang et al. (2024); Li et al. (2023b); Yu et al. (2024b); Li et al. (2025a); Kuang et al. (2024), recent studies Chu et al. (2025); Ye et al. (2025c, b, 2024); Hu et al. (2024); Chen et al. (2024); Su et al. (2025); Zou et al. (2025); Xu et al. (2025); Li et al. (2025b, 2024b); Zhang et al. (2025b); Li et al. (2024a) have generated interest in employing LLMs for text assessment on various NLP tasks. Building on this idea, we use Llama-2-7B Touvron et al. (2023) as a scorer to determine edit weights. The prompt for edit weighting is presented in Figure 3. We set the temperature to 0.1 to ensure consistent and certain results. We instruct the LLM to evaluate each edit individually to prevent interference from other grammatical errors. Edit weights vary from 1 to 5, with higher values representing a greater need for correction. We do not specify the types of edits to the LLM; instead, we allow the LLM to directly evaluate the importance of edits through its inherent language understanding abilities. An input is composed of an uncorrected sentence and a certain edit.
Appendix B Details about GEC Meta-Evaluation
B.1 Human Rankings
GJG15 ranking.
Grundkiewicz et al. (2015) propose the first large-scale human judgement dataset for 12 participating systems of the CoNLL-2014 shared task. In this assessment, 8 native speakers are asked to rank the systems’ outputs from best to worst. Two system ranking lists are generated using Expected Wins (EW) and TrueSkill (TS), respectively.
SEEDA ranking.
Kobayashi et al. (2024b) identify several limitations of the GJG15 ranking dataset, and propose a new human ranking dataset called SEEDA. SEEDA consists of corrections with human ratings along two different granularities: edit-based and sentence-based, covering 12 state-of-the-art systems, including large language models (LLMs), and two human corrections with different focuses. Three native English speakers participate in the annotation process. Similar to Grundkiewicz et al. (2015), the overall human rankings are derived from TrueSkill (TS) and Expected Wins (EW) based on pairwise judgments.
B.2 Ranking Algorithms
Our employed human judgments are originally pairwise comparisons, i.e., humans choose the better of two available system outputs. The overall rankings are derived by using ranking algorithms, including Expected Wins (EW) and TrueSkill (TS).
Expected Wins (EW)
EW Bojar et al. (2013) is a derived ranking metric that quantifies the theoretical number of wins a participant is expected to achieve against a defined set of opponents. It is calculated by summing the probability of winning against each opponent, where these probabilities are typically derived from an existing skill rating system. EW provides a single aggregate score for ranking, useful for pre-match seeding or assessing theoretical group performance.
TrueSkill (TS)
TS Sakaguchi et al. (2014) is a Bayesian skill rating system developed by Microsoft Research. Unlike simpler systems, TS models a participant’s skill as a probability distribution (), where represents the estimated skill level and quantifies the uncertainty in that estimate. Upon match outcomes, TS updates these distributions using Bayesian inference, allowing for rapid adjustments and robust ranking. A key advantage is its inherent support for multi-player or team-based matches and the explicit handling of draws. Participants are typically ranked by a conservative estimate of their skill, such as , which accounts for confidence.
| # Sentence (Length) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) |
| # Reference (Length) | 2,624 (22.8) | 13,120 (22.9) | 2,624 (23.2) | 2,624 (22.8) | 2,624 (23.0) | 2,624 (22.2) |
| # Edit (Length) | 5,937 (1.0) | 36,677 (1.0) | 4,500 (1.0) | 8,373 (1.1) | 4,964 (0.9) | 11,033 (1.2) |
| # Unchanged Chunk (Length) | 11,174 (4.8) | 93,496 (2.5) | 8,887 (6.3) | 12,823 (3.8) | 10,748 (5.1) | 14,086 (2.9) |
| # Corrected/Dummy Chunk (Length) | 4,994 (1.3) | 26,948 (2.4) | 3,963 (1.2) | 6,305 (1.7) | 4,221 (1.2) | 6,892 (2.6) |
| EW | TS | EW | TS | EW | TS | EW | TS | EW | TS | EW | TS | |||
| SentGLEU | 0.784 | 0.828 | 0.756 | 0.826 | 0.624 | 0.581 | 0.785 | 0.846 | 0.218 | 0.142 | 0.778 | 0.848 | 0.668 ( 0.120) | |
| 0.720 | 0.775 | 0.769 | 0.824 | 0.599 | 0.593 | 0.791 | 0.846 | 0.220 | 0.170 | 0.768 | 0.846 | 0.660 ( 0.131) | ||
[…]
| 0.742 | 0.747 | 0.786 | 0.830 | 0.626 | 0.588 | 0.813 | 0.764 | -0.003 | -0.137 | 0.830 | 0.857 | 0.620 ( 0.179) | ||
[…]
| 0.824 | 0.808 | 0.835 | 0.874 | 0.451 | 0.412 | 0.824 | 0.863 | -0.181 | -0.247 | 0.791 | 0.846 | 0.592 ( 0.233) | ||
[…]
| 0.725 | 0.758 | 0.209 | 0.143 | 0.368 | 0.335 | 0.808 | 0.846 | -0.167 | -0.247 | 0.762 | 0.825 | 0.447 ( 0.335) | ||
[…]
| 0.714 | 0.725 | 0.681 | 0.747 | 0.846 | 0.852 | 0.824 | 0.901 | 0.368 | 0.352 | 0.720 | 0.791 | 0.710 ( 0.091) | ||
[…]
| 0.709 | 0.720 | 0.681 | 0.747 | 0.819 | 0.835 | 0.857 | 0.923 | 0.423 | 0.401 | 0.654 | 0.720 | 0.707 ( 0.086) | ||
[…]
| 0.907 | 0.912 | 0.808 | 0.863 | 0.808 | 0.813 | 0.885 | 0.945 | 0.527 | 0.505 | 0.896 | 0.940 | 0.817 ( 0.051) | ||
[…]
| 0.868 | 0.879 | 0.753 | 0.824 | 0.808 | 0.841 | 0.879 | 0.956 | 0.544 | 0.527 | 0.835 | 0.923 | 0.803 ( 0.052) | ||
B.3 Statistics of Reference Datasets
B.4 Baseline Metrics
ERRANT and SentERRANT Bryant et al. (2017) are among the most widely recognized in grammatical error correction. They enhance the accuracy of edit extraction by employing a linguistically refined version of the Damerau-Levenshtein algorithm.
CLEME and SentCLEME Ye et al. (2023c) are proposed to provide unbiased scores for multi-reference evaluation. Furthermore, the authors present the correction independence assumption, enabling CLEME to function under either the traditional correction dependence or correction independence assumptions.
| Chunk 1 | Chunk 2 | Chunk 3 | Chunk 4 | Chunk 5 | Chunk 6 | |
| Reference | Does one | who | suffers | from this disease keep it a secret | or inform | their relatives ? |
| Hypothesis | Do one (5) | who | suffer (5) | from this disease keep it a secret | to inform (1) | their relatives ? |
| Source | When we are | diagonosed out | with certain genetic | disease | , should we disclose | this result | to | our | relatives ? |
| Ref. | When we are | diagnosed | with certain genetic | diseases | , should we disclose | this result | to | our | relatives ? |
| Hyp. | When we are | diagnosed out (5) | with certain genetic | diseases (5) | , should we disclose | the results (4) | to | their (5) | relatives ? |
For the evaluation on SEEDA, we add extra evaluation metrics following the evaluation methods reported in Kobayashi et al. (2024b):
SOME Yoshimura et al. (2020): optimizes human evaluation by fine-tuning BERT separately for criteria such as grammaticality, fluency, and meaning preservation.
B.5 Details of Determining Trade-off Factors
A cross-validation approach was employed on the six reference sets of GJG15 to determine the optimal set. Five of the six reference sets were selected, and an exhaustive exploration of all trade-off factors was conducted. The candidate factors were evaluated at intervals determined by a grid value of 0.05. The optimal factors were then identified and applied to the remaining reference set, yielding resultant corrections. We reiterated this process six times to ascertain the final set of trade-off factors, which exhibited the highest average correction for the remaining reference sets.
Appendix C Extra Results
C.1 Results of Full References
The results without excluding unchanged reference sentences are presented in Table 8. We observe an obvious performance reduction in traditional metrics, especially in NE-Minimal, which contains numerous under-corrections due to annotation by non-experts under the minimal editing guideline. We remove 470 unchanged references in E-Minimal and 612 unchanged references in NE-Minimal. In particular, SentERRANT, SentCLEME-dep, and SentCLEME-ind exhibit negative correlations in NE-Minimal, revealing their lack of robustness. Many metrics also undergo a significant decrease in E-Minimal except CLEME2.0. In the case of E-Minimal, many metrics also show a marked decline, except for CLEME2.0. Our approach achieves the highest or comparable correlations across all reference sets, underscoring its robustness.
C.2 Case Study of LLM-based Weighting
In Table 9, we report instances of CLEME2.0 using LLM-based weighting. We notice distinct preferences when comparing similarity-based and LLM-based weighting methods. In the first example, Llama-2 attributes significant weights to Chunks 1 and 2, highlighting key grammatical mistakes. Conversely, it assigns a minor weight to Chunk 5 due to its imperfect modification. The second example shows Llama-2 attributing substantial weights to all chunks. Specifically, for Chunk 2, the hypothesis fails to remove the redundant “out," emphasizing the under-correction issue. Chunks 6 and 8 display excessive corrections, altering the sentence’s original intent and thus indicating considerable over-correction. Generally, Llama-2 tends to ascribe either very high or very low weights to modifications. We speculate it is due to the small scale of the LLM we adopt, impairing its ability to distinguish grammatical errors with varying levels.
| Metric | EW | TS | ||
| CLEME-dep-unw | 0.638 | 0.654 | 0.681 | 0.709 |
| CLEME-ind-unw | 0.640 | 0.648 | 0.680 | 0.698 |
| CLEME-dep-len | 0.700 | 0.665 | 0.691 | 0.742 |
| CLEME-ind-len | 0.649 | 0.709 | 0.691 | 0.731 |
| CLEME-dep-sim | 0.655 | 0.764 | 0.698 | 0.797 |
| CLEME-ind-sim | 0.641 | 0.720 | 0.687 | 0.747 |
| SentCLEME-dep-unw | 0.853 | 0.687 | 0.805 | 0.604 |
| SentCLEME-ind-unw | 0.790 | 0.275 | 0.722 | 0.181 |
| SentCLEME-dep-len | 0.876 | 0.824 | 0.844 | 0.808 |
| SentCLEME-ind-len | 0.868 | 0.725 | 0.857 | 0.758 |
| SentCLEME-dep-sim | 0.888 | 0.692 | 0.844 | 0.648 |
| SentCLEME-ind-sim | 0.843 | 0.500 | 0.786 | 0.434 |
C.3 Extra Results of CLEME with Similarity Weighting
We additionally investigate the application of similarity-based weighting to CLEME Ye et al. (2023c) and present the results on CoNLL2014 in Table 10. We find that similarity-based weighting is superior to length-based weighting for corpus-level CLEME, while the trend is reversed for SentCLEME, and both are better than the unweighted setting. Moreover, it should be noted that no matter the weighting strategy employed, CLEME consistently underperforms compared to CLEME2.0. This is attributed to the fundamental disparities in design and scoring frameworks between the versions. CLEME2.0 was crafted to incorporate these sophisticated weighting techniques, allowing it to better distinguish between diverse error types and deliver a more thorough and refined performance assessment.
@@ C.3 Extra Results of CLEME with Similarity Weighting
C.4 Average Correlations.